# GaussianTalker: Real-Time High-Fidelity Talking Head Synthesis with Audio-Driven 3D Gaussian Splatting

### v1

@@ Abstract.
We propose GaussianTalker, a novel framework for real-time generation of pose-controllable talking heads. It leverages the fast rendering capabilities of 3D Gaussian Splatting (3DGS) while addressing the challenges of directly controlling 3DGS with speech audio. GaussianTalker constructs a canonical 3DGS representation of the head and deforms it in sync with the audio. A key insight is to encode the 3D Gaussian attributes into a shared implicit feature representation, where it is merged with audio features to manipulate each Gaussian attribute. This design exploits the spatial-aware features and enforces interactions between neighboring points. The feature embeddings are then fed to a spatial-audio attention module, which predicts frame-wise offsets for the attributes of each Gaussian. It is more stable than previous concatenation or multiplication approaches for manipulating the numerous Gaussians and their intricate parameters. Experimental results showcase GaussianTalker’s superiority in facial fidelity, lip synchronization accuracy, and rendering speed compared to previous methods. Specifically, GaussianTalker achieves a remarkable rendering speed of 120 FPS, surpassing previous benchmarks.
@@
ccs: Computing methodologies Reconstructionccs: Information systems Multimedia content creationccs: Computing methodologies 3D imaging
Figure 1. Fidelity and inference time comparison between existing 3D talking face synthesis models (Guo et al., 2021; Tang et al., 2022; Li et al., 2023) and ours. Our method, GaussianTalker, achieves on par with or better results at much higher FPS. Note that we also include GaussianTalker, a more efficient and faster variant. Size of each bubble represents the inference time of each method.
@@ 2.3. Facial animation with 3DGS
Previous methods for facial reconstrution and animation primarily relied on 3D Morphable Models(3DMM) (Grassal et al., 2022; Khakhulin et al., 2022) or utilized neural implicit representations (Zheng et al., 2022; Athar et al., 2022; Gao et al., 2022). Recent approaches (Qian et al., 2023; Wang et al., 2024; Chen et al., 2023; Dhamo et al., 2023) have shifted towards adopting the 3DGS representation, aiming to leverage the benefits of rapid training and rendering while still achieving competitive levels of photorealism. GaussianAvatars (Qian et al., 2023) reconstructed head avatars by rigging 3D Gaussians on FLAME (Li et al., 2017) mesh. MonoGaussianAvatar (Chen et al., 2023) learned explicit head avatars by shifting the mean position of 3D Gaussians from canonical to deformed space using Linear Blend Skinning (LBS) and simultaneously adjusts other Gaussian parameters through a deformation field. GaussianHead (Wang et al., 2024) adopted a motion deformation field to adapt to facial movements while preserving head geometry and separately utilized a tri-plane to retain the appearance information of individual 3D Gaussians. However, the aforementioned methods tend to depend on parametric models for facial animation. In contrast to previous works, our audio-driven method is not only free from the need for data beyond the speech sequence for facial reenactment but also is readily applicable to novel audio.
@@ 4.2.1. Triplane representation for 3D Gaussian
Figure 3. Visualization of the triplane feature grids. The sequence displays a rendered image, followed by its orthographic projected embeddings: frontal (xy), overhead (yz), and side (zx) views. The embeddings are visualized by reducing its dimenstion to 3 using Principal Component Analysis.
@@ 4.3. Learning audio-driven deformation of 3D Gaussians
4.3.1. Spatial-audio cross-attention
Previous approaches to implement region-aware audio, like ER-NeRF (Li et al., 2023), simply adjust the weights for the audio features at each 3D point through elementwise multiplication. However, it encounters a challenge in that, regardless of the diverse audio inputs in a dynamic scene, a particular static 3D point consistently maintains the same audio weight. This fails to acknowledge that a fixed 3D coordinate may not consistently correspond to the same facial region as the scene progresses. To address this issue and enhance the extraction of spatial-audio features, we introduce spatial-audio cross-attention module, a cross-attention mechanism that merges spatial feature embedding $f(\mu_{c})$ of the canonical 3D Gaussians with subsequent audio features, capturing how the input speech audio affects the movement of the 3D Gaussians. The spatial-audio cross-attention module comprises $L$ sets of cross-attention layer $\mathcal{T}_{CA}(\cdot)$ and feed-forward layer $FFN(\cdot)$, each interconnected with skip connections. The module is formulated as:
@@ 4.3.3. Audio-conditioned deformation of 3D Gaussian
Table 1. Quantitative comparison under the self-driven setting. The top, second-best, and third-best results are shown in red, orange, and yellow, respectively.
@@ 5.1. Experimental Settings
Figure 5. Comparative visualization of lip synchronization across different audio-visual models. The sequence depicts the lip shape conforming to specific phonemes in the spoken words ’country’, ’of’, ’crime’, ’we’, ’up’, ’especially’, ’like’, with the last frame showing a closed mouth (’mute’).
@@ 5.2. Quantitative Evaluation
Following previous works (Tang et al., 2022; Li et al., 2023), our comparisons are structured into two distinct settings: self-driven and cross-driven. In the self-driven setting, we evaluate the accuracy of head reconstruction for a particular identity using the test subset. We employ several reconstruction metrics including peak signal-to-noise ratio (PSNR), structural similarity index measure (SSIM), and learned perceptual image patch similarity (LPIPS). Notably, these metrics are exclusively measured on the facial region. We also measure realism of the reconstructed face using Fréchet Inception Distance (FID(Heusel et al., 2017) and identity preservation of the animated video using Cosine Similarity of Identity Embedding (CSIM(Huang et al., 2020).
@@ 5.4. Ablation Study
Table 3. Ablation study results comparing various attribute configurations for embedding canonical 3D Gaussian attributes.
@@ 6. Conclusion
In this work, we have proposed GaussianTalker, a novel framework for real-time pose-controllable 3D talking head synthesis, leveraging the 3D Gaussians for the head representation. Our method enables precise control over Gaussian primitives by conditioning features extracted from a multi-resolution triplane. Additionally, the integration of a spatial-audio cross-attention module facilitates the dynamic deformation of facial regions, allowing for nuanced adjustments based on audio cues and enhancing verbal motion disentanglement. Our method is distinguished from prior NeRF-based methods by its superior inference speed and high-fidelity results for out-of-domain audio tracks. The efficacy of our approach is validated by quantitative and qualitative analyses. We look forward to enriched user experiences, particularly in video game development, where real-time rendering capabilities of GaussianTalker promise to enhance interactive digital environments.
@@ A.2. Hyperparameter Configuration
While our spatial-audio cross-attention module primarily employs $L=2$ cross-attention layers, our modified GaussianTalker with $L=1$ can achieve comparable results with even faster inference speeds.
@@ B.1. Splatting on the background image
As a solution, we opted to generate faces against GT backgrounds instead. This approach allowed for the accurate learning of Gaussian presence boundaries by distributing loss across the entire image. Similar to preprocessing techniques employed in previous NeRF-based works (Guo et al., 2021; Tang et al., 2022; Li et al., 2023), we interpolated the human form from the background image to create an image with the person removed. Subsequently, faces were directly rendered using Gaussian methods, enabling comparisons with GT videos. By adopting this strategy, our GaussianTalker is trained without the need for facial mask, facilitating the faithful representation of intricate details such as hair.
@@ B.2. Visualization of Attention
In our spatial-audio cross-attention module, the computation of the attention score is formalized by the following equation:
(19) $(\mathrm{A}_{n})^{l}=\frac{\mathrm{softmax}(qk^{\intercal}_{n})^{l}}{\sqrt{d_{k}}},$
where $l$ denotes the index of $\{a_{n},e_{n},\upsilon_{n},\emptyset\}$ and $(\mathrm{A}_{n})^{l}$ corresponds to its calculated attention score. $\mathrm{A}_{n}$ denotes the concatenation of all $(\mathrm{A}_{n})^{l}$, resulting in a shape of ${{B}\times{H}\times{N}\times{d_{k}}}$, which respectively indicate batch size, number of heads, number of Gaussians, and number of features per Gaussian.
For each attention score $(\mathrm{A}_{n})^{l}$, we visualize the attention by assigning the score to RGB values. Thereby we obtain attention visualization colors $c_{att}$ for each Gaussian. The overall visualization of attention is then calculated such as:
where $c_{i}$ represents the color associated with each Gaussian, determined by $c_{att}$ along the view direction. ${\alpha}_{i}^{\prime}$ is derived from the multiplication of the opacity $\alpha$ of the 3D Gaussian and its projected covariance $\Sigma^{\prime}$. This mathematical formulation allows us to visually interpret the model’s focus within the generated representations, effectively highlighting the areas of greatest feature impact.
@@ Appendix E Supplementary Video
To comprehensively visualize the efficacy of our proposed method in the domain of talking facial video synthesis, we prepared a supplementary video. This video encompasses the results and analysis of our experiments presented in the main paper and the supplementary document. We showcase talking head videos generated under both the self-driven and cross-driven settings and compare them with previous NeRF-based works (Guo et al., 2021; Tang et al., 2022; Li et al., 2023). We also demonstrate the effectiveness of our spatial-audio cross attention module by showing how the attention scores of each condition evolve as the scene progresses. Lastly, the video includes a set of ablation studies that systematically examine the impact of each component of our proposed method.

### v2

@@ Abstract.
We propose GaussianTalker, a novel framework for real-time generation of pose-controllable talking heads. It leverages the fast rendering capabilities of 3D Gaussian Splatting (3DGS) while addressing the challenges of directly controlling 3DGS with speech audio. GaussianTalker constructs a canonical 3DGS representation of the head and deforms it in sync with the audio. A key insight is to encode the 3D Gaussian attributes into a shared implicit feature representation, where it is merged with audio features to manipulate each Gaussian attribute. This design exploits the spatial-aware features and enforces interactions between neighboring points. The feature embeddings are then fed to a spatial-audio attention module, which predicts frame-wise offsets for the attributes of each Gaussian. It is more stable than previous concatenation or multiplication approaches for manipulating the numerous Gaussians and their intricate parameters. Experimental results showcase GaussianTalker’s superiority in facial fidelity, lip synchronization accuracy, and rendering speed compared to previous methods. Specifically, GaussianTalker achieves a remarkable rendering speed up to 120 FPS, surpassing previous benchmarks. The code is made public in https://ku-cvlab.github.io/GaussianTalker/.
@@
copyright: noneconference: Make sure to enter the correct conference title from your rights confirmation email; ; ccs: Computing methodologies Reconstructionccs: Information systems Multimedia content creationccs: Computing methodologies 3D imagingfootnotetext: * Contributed equally to this research.footnotetext: $\dagger$ Corresponding author.
Figure 1. Fidelity, lip synchronization and inference time comparison between existing 3D talking face synthesis models (Guo et al., 2021; Tang et al., 2022; Li et al., 2023) and ours. Our method, GaussianTalker, achieves superior performance at much higher FPS. Note that we also include GaussianTalker, a more efficient and faster variant. Size of each bubble represents the inference time per frame of each method.
@@ 2.3. Facial animation with 3DGS
Previous methods for facial reconstruction and animation primarily relied on 3D Morphable Models(3DMM) (Grassal et al., 2022; Khakhulin et al., 2022) or utilized neural implicit representations (Zheng et al., 2022; Athar et al., 2022; Gao et al., 2022). Recent approaches (Qian et al., 2023; Wang et al., 2024; Chen et al., 2023; Dhamo et al., 2023) have shifted towards adopting the 3DGS representation, aiming to leverage the benefits of rapid training and rendering while still achieving competitive levels of photorealism. GaussianAvatars (Qian et al., 2023) reconstructed head avatars by rigging 3D Gaussians on FLAME (Li et al., 2017) mesh. MonoGaussianAvatar (Chen et al., 2023) learned explicit head avatars by shifting the mean position of 3D Gaussians from canonical to deformed space using Linear Blend Skinning (LBS) and simultaneously adjusts other Gaussian parameters through a deformation field. GaussianHead (Wang et al., 2024) adopted a motion deformation field to adapt to facial movements while preserving head geometry and separately utilized a tri-plane to retain the appearance information of individual 3D Gaussians. However, the aforementioned methods tend to depend on parametric models for facial animation. In contrast to previous works, our audio-driven method is not only free from the need for data beyond the speech sequence for facial reenactment but also is readily applicable to novel audio.
@@ 4.2.1. Triplane representation for 3D Gaussian
Figure 3. Visualization of the triplane feature grids. The sequence displays a rendered image, followed by its orthographic projected embeddings: frontal (xy), overhead (yz), and side (zx) views. The embeddings are visualized by reducing its dimension to 3 using PCA.
@@ 4.3. Learning audio-driven deformation of 3D Gaussians
4.3.1. Spatial-audio cross attention
Previous approaches to implement region-aware audio, like ER-NeRF (Li et al., 2023), simply adjust the weights for the audio features at each 3D point through elementwise multiplication. However, it encounters a challenge in that, regardless of the diverse audio inputs in a dynamic scene, a particular static 3D point consistently maintains the same audio weight. This fails to acknowledge that a fixed 3D coordinate may not consistently correspond to the same facial region as the scene progresses. To address this issue and enhance the extraction of spatial-audio features, we introduce spatial-audio attention module, a cross-attention mechanism that merges spatial feature embedding $f(\mu_{c})$ of the canonical 3D Gaussians with subsequent audio features, capturing how the input speech audio affects the movement of the 3D Gaussians. The spatial-audio attention module comprises $L$ sets of cross-attention layer $\mathcal{T}_{CA}(\cdot)$ and feed-forward layer $FFN(\cdot)$, each interconnected with skip connections. The module is formulated as:
@@ 4.3.3. Audio-conditioned deformation of 3D Gaussian
Table 1. Quantitative comparison under the self-driven setting. The top, second-best, and third-best results are shown in red, orange, and yellow, respectively. Our GaussianTalker and the lightweight version GaussianTalker achieves performance on par with or better results at significantly lower inference time.
@@ 5.1. Experimental Settings
Figure 5. Comparative visualization of lip synchronization across different audio-visual models. The sequence depicts the lip shape conforming to specific phonemes in the spoken words ‘country’, ‘of’, ‘crime’, ‘we’, ‘up’, ‘especially’, ‘like’, with the last frame showing a closed mouth (‘mute’).
@@ 5.2. Quantitative Evaluation
Following previous NeRF-based works (Tang et al., 2022; Li et al., 2023), our comparisons are structured into two distinct settings: self-driven and cross-driven. In the self-driven setting, we evaluate the accuracy of head reconstruction for a particular identity using the test subset. We employ several reconstruction metrics including peak signal-to-noise ratio (PSNR), structural similarity index measure (SSIM), and learned perceptual image patch similarity (LPIPS). Notably, these metrics are exclusively measured on the facial region. We also measure realism of the reconstructed face using Fréchet Inception Distance (FID(Heusel et al., 2017) and identity preservation of the animated video using Cosine Similarity of Identity Embedding (CSIM(Huang et al., 2020).
@@ 5.4. Ablation Study
5.4.1. Attribute conditions for triplane
Table 3. Ablation study results comparing various attribute configurations to be conditioned by $f(\mu_{c})$ for the canonical 3D head.
@@ 6. Conclusion
In this work, we have proposed GaussianTalker, a novel framework for real-time pose-controllable 3D talking head synthesis, leveraging the 3D Gaussians for the head representation. Our method enables precise control over Gaussian primitives by conditioning features extracted from a multi-resolution triplane. Additionally, the integration of a spatial-audio attention module facilitates the dynamic deformation of facial regions, allowing for nuanced adjustments based on audio cues and enhancing verbal motion disentanglement. Our method is distinguished from prior NeRF-based methods by its superior inference speed and high-fidelity results for out-of-domain audio tracks. The efficacy of our approach is validated by quantitative and qualitative analyses. We look forward to enriched user experiences, particularly in video game development, where real-time rendering capabilities of GaussianTalker promise to enhance interactive digital environments.
@@ A.2. Hyperparameter Configuration
While our spatial-audio attention module primarily employs $L=2$ cross-attention layers, our modified GaussianTalker with $L=1$ can achieve comparable results with even faster inference speeds.
@@ B.1. Splatting on the background image
As a solution, we opted to generate faces GT backgrounds instead. This approach allowed for the accurate learning of Gaussian presence boundaries by distributing loss across the entire image. Similar to preprocessing techniques employed in previous NeRF-based works (Guo et al., 2021; Tang et al., 2022; Li et al., 2023), we interpolated the human form from the background image to create an image with the person removed. Subsequently, faces were directly rendered using Gaussian methods, enabling comparisons with GT videos. By adopting this strategy, our GaussianTalker is trained without the need for facial mask, facilitating the faithful representation of intricate details such as hair.
@@ B.2. Visualization of Attention
In our spatial-audio attention module, the computation of the attention score is formalized by the following equation:
(19) $\mathrm{A}_{n}^{l}=\frac{\mathrm{softmax}(qk^{\intercal}_{n})^{l}}{\sqrt{d_{k}}},$
where $l$ denotes the index of $\{a_{n},e_{n},\upsilon_{n},\emptyset\}$ and $\mathrm{A}_{n}^{l}$ corresponds to its calculated attention score. $\mathrm{A}_{n}$ denotes the concatenation of all $\mathrm{A}_{n}^{l}$, resulting in a shape of ${{B}\times{H}\times{N}\times{d_{k}}}$, which respectively indicate batch size, number of heads, number of Gaussians, and number of features per Gaussian.
For each attention score $\mathrm{A}_{n}^{l}$, we visualize the attention by assigning the score to RGB values. Thereby we obtain attention visualization colors $c_{att}$ for each Gaussian. The overall visualization of attention is then calculated with the typical 3DGS rendering process following (5). This formulation allows us to visually interpret the model’s focus within the generated representations, effectively highlighting the areas of greatest feature impact.
@@ Appendix E Supplementary Video
To comprehensively visualize the efficacy of our proposed method in the domain of talking facial video synthesis, we prepared a supplementary video. This video encompasses the results and analysis of our experiments presented in the main paper and the supplementary document. We showcase talking head videos generated under both the self-driven and cross-driven settings and compare them with previous NeRF-based works (Guo et al., 2021; Tang et al., 2022; Li et al., 2023). We also demonstrate the effectiveness of our spatial-audio attention module by showing how the attention scores of each condition evolve as the scene progresses. Lastly, the video includes a set of ablation studies that systematically examine the impact of each component of our proposed method.